Cougar Plastics Company has been operating for three years. At December 31 of last year, the accounting records reflected the following:

Cash $ 20,000
Accounts payable $ 19,000
Investments (short-term) 3,500
Accrued liabilities payable 2,000
Accounts receivable 4,400
Notes payable (short-term) 5,200
Inventory 30,000
Notes payable (long-term) 49,000
Notes receivable (long-term) 1,400
Common stock 9,300
Equipment 43,000
Additional paid-in capital 83,700
Factory building 101,000
Retained earnings 38,600
Intangibles 3,500

Summarized activities during the year:
Purchased short-term investments for $7,800 cash.
Lent $5,900 to a supplier who signed a two-year note.
Purchased equipment that cost $23,000; paid $4,500 cash and signed a one-year note for the balance.
Hired a new president at the end of the year. Contract was for $88,000 per year plus options to purchase company stock at a set price based on company performance.
Issued an additional 1,300 shares of $0.50 par value common stock for $19,000 cash.
Borrowed $15,000 cash from a local bank, payable in three months.
Purchased a patent (an intangible asset) for $1,200 cash.
Built an addition to the factory for $24,000; paid $8,000 in cash and signed a three-year note for the balance.
Returned defective equipment to the manufacturer, receiving a cash refund of $3,400.

1-4:

Post the current year transactions to T-accounts for each of the accounts on the balance sheet (Cash, Investments, Accounts Receivable, Inventory, Notes receivable long term, equipment, factory building, intangibles, accounts payable, accrued liabilities payable, notes payable short term, long term notes payable, common stock, additional paid-in capital, retained earnings), Prepare a trial balance at December 31 of the current year, & Compute the current ratio for the current year -- round to 2 decimal places.
Answer #1

ACCOUNTS AS ON 31ST DECEMBER

PARTICULARS AMT(IN $) PARTICULARS AMT(IN $)
1. CASH ACCOUNT
TO OPENING 20000 BY ST INVESTMENT 7800
TO ISSUE OF COMMON STOCK 19000 BY LENT (NOTE RECEIVABLES LONG TERM) 5900
TO NOTE PAYABLE SHORT TERM 15000 BY EQUIPMENT PURCHASE 4500
TO EQUIPMENT RETURNED 3400 BY PATENT PURCHASE 1200
BY FACTORY ADDITION 8000

BY PAYMENT TO BANK FOR BORROWINGS (ST NP)

15000
BY CLOSING 15000
57400 57400
2.ACCOUNT PAYABLE
BY OPENING 19000
TO CLOSING 19000
19000 19000
3. SHORT TERM INVESTMENT
TO OPENING 3500
TO CASH 7800 BY CLOSING 11300
11300 11300
4. ACCRUED LIABILITIES PAYABLE
BY OPENING 2000
TO CLOSING 2000
2000 2000
5.ACCOUNT RECEIVABLES
TO OPENING 4400 BY CLOSING 4400
4400 4400
6. INVENTORY
TO OPENING 30000 BY CLOSING 30000
30000 30000
7. NOTE PAYABLE SHORT TERM
TO CASH ( PAYMENT OF BORROWING) 15000 BY OPENING 5200
TO EQUIPMENT PURCHASE 18500
TO CLOSING 23700 TO CASH 15000
38700 38700
8. NOTE PAYABLE LONG TERM
BY OPENING 49000
TO CLOSING 65000 BY FACTORY 16000
65000 65000
9. NOTE RECEIVABLES LONG TERM
TO OPENING 1400
TO CASH 5900 BY CLOSING 7300
7300 7300
10. COMMON STOCK
BY OPENING 9300
BY PRESIDENT (W.N) 3011
TO CLOSING 12961 BY CASH (1300 SHARES @ .50 PAR) 650
12961 12961
11. ADDITIONAL PAID-UP CAPITAL
BY OPENING 83700
TO CLOSING 83700
83700 83700
12. RETAINED EARNING
BY OPENING 38600
TO PRESIDENT (P&L) 88000 BY CASH (19000-650) (SEC. PREMIUM) 18350
TO CLOSING 53939 BY PRESIDENT (SEC. PREMIUM) (W.N) 84989
141939 141939
13. EQUIPMENT
TO OPENING 43000
TO PURCHASE 23000 BY RETURNED (CASH)

3400

BY CLOSING 62600
66000 66000
14. FACTORY BUILDING
TO OPENING 101000
TO PURCHASE 24000 BY CLOSING 125000
125000 125000
15. INTANGIBLE ASSET`
TO OPENING 3500
TO CASH (PATENT PURCHASE) 1200 BY CLOSING 4700
4700 4700
16. PRESIDENT
TO COMMON STOCK 3011 BY SALARIES / REMUNERATION (P & L) 88000
TO SEC. PREMIUM 84989
88000 88000

NOTE: WE ASSUMED THAT BORROWING FROM BANK $ 15000 FOR THREE MONTH ON THE BASIS OF SHORT TERM NOTE PAYABLE IS PAID IN CURRENT YEAR.

WORKING NOTE:

1. FACE VALUE OF SHARE = $ 0.50PER SHARE

NO. OF ADDITION SHARE ISSUED = 1300

FOR 13000 CASH

SHARE CAPITAL = 1300 @ 0.50 = $ 650

SECURITY PREMIUM = 19000-650 = 18350.

2. CONTRACT WITH PRESIDENT FOR $ 88000. WE ASSUMED THAT IS FOR REMUNERATION / SALARY WHICH IS PAID BY ISSUE COMMON STOCK.

NO. OF COMMON STOCK ISSUED = 88000 / 19000 X 1300 = 6021 SHARES (APPROX)

SHARE CAPITAL = 6021 @ 0.50 = $ 3011 (APPROX)

SECURITY PREMIUM = 88000 - 3011 = 84989.

OR

IN RATIO

SHARE CAPITAL = 88000 x 650 /19000 = 3011 (APPROX)

SEC. PREMIUM = 88000 X 18350 / 19000 = 84989 (APPROX).

TRIAL BALANCE AS ON 31 ST DECEMBER DR. CR.
PARTICULARS AMT (IN $) AMT (IN $)
COMMON STOCK 12961
RETAINED EARNING 53939
ADDITIONAL PAID-UP CAPITAL 83700
FACTORY BUILDING 125000
EQUIPMENT 62600
INTANGIBLE ASSET 4700
LONG TERM NOTE RECEIVABLES / PAYABLES 7300 65000
ACCOUNT RECEIVABLES / PAYABLES 4400 19000
SHORT TERM INVESTMENT 11300
CASH 15000
ACCRUED LIABILITIES PAYABLE 2000
SHORT TERM NOTE PAYABLE 23700
OPENING INVENTORY 30000
TOTAL 260300 260300

NOTE : CLOSING STOCK = 30000.

CALCULATION OF CURRENT RATIO:

CURRENT RATIO = CURRENT ASSETS / CURRENT LIABILITIES

= 60700 / 44700

= 1.36 TIMES

CURRENT ASSETS = CLOSING INVENTORY + CASH + ST INVESTMENT + ACCOUNT RECEIVABLES

= 30000 + 15000 + 11300 + 4400

= 60700

CURRENT LIABILITIES = ACCOUNT PAYABLE + ACCRUED LIABILITIES + ST NOTE PAYABLE

= 19000 + 2000 + 23700

= 44700